OpenStudy (anonymous):

what is the slope of the line that passes through the points (3,5) and (-2,2)?

OpenStudy (anonymous):

slope of a line is (y2-y1)/(x2 - x1)

OpenStudy (anonymous):

1/7

OpenStudy (anonymous):

thats the slope

OpenStudy (anonymous):

nope its 5/3 :p

OpenStudy (anonymous):

its 5/3

OpenStudy (radar):

\[m=\Delta y/\Delta x\]

OpenStudy (radar):

m=(2-5)/(-2-3)=-3/-5=3/5

OpenStudy (radar):

your points are x=3, and y=5 and x=-2, and y=2

OpenStudy (radar):

The formula provided by damdam is a valid one, just do the math.